# Benchmarking Algorithms for Federated Domain Generalization

### v1

@@ 4 Benchmark Experimental Results
Table 3: Test accuracy on PACS and IWildCam dataset with held-out-domain validation where FedAvg-ERM is the simple baseline (B). “-” means the method is not applicable in that context. Bold is for best and italics is for second best in each column.
@@
FedAvg-ERM $0.8379$ $0.8766$ $0.9144$ $\mathit{0.9377}$ $0.2727$ $0.1762$ $0.1241$ $0.0707$
DG Adapted IRM $0.9180$ $0.8156$ $0.5449$ $0.1669$ $0.0718$ $0.0903$
Fish $0.9030$ $0.8497$ $0.6311$ $0.2103$ $0.0781$ $0.0977$
Mixup $0.8635$ $0.7653$ $0.5551$ $0.1323$ $0.0264$ $0.0370$
MMD $\mathit{0.9186}$ $0.8150$ $0.6341$ $\mathit{0.3134}$ $0.0718$ $0.0924$
DeepCoral $\bf 0.9216$ $0.8150$ $0.5515$ $\mathbf{0.3270}$ $0.0718$ $0.0926$
GroupDRO $0.9060$ $\bf 0.9395$ $\bf 0.9437$ $0.2531$ $0.0530$ $0.0908$
FL FedProx - $0.8826$ $0.8407$ $0.8892$ $\mathbf{0.2220}$ $\mathbf{0.1770}$ $\mathbf{0.1535}$
Scaffold - $0.8215$ $0.8796$ $0.9239$ $\mathit{0.2098}$ $\mathit{0.1532}$ $\mathit{0.1467}$
FDG FedDG - $\mathit{0.9234}$ $\mathit{0.9275}$ $\mathbf{0.9521}$ $0.1445$ $0.1191$ $0.0503$
FedADG - $0.1892$ $0.0592$ $0.0598$ $0.0049$ $0.0049$ $0.0049$
FedSR - $0.1246$ $0.1263$ $0.1257$ $0.0056$ $0.0056$ $0.0056$
FedGMA - $0.8820$ $0.8467$ $0.8446$ $0.0106$ $0.0957$ $0.0106$
Table 4: Test accuracy on CivilComments and Py150 datasets with held-out-domain validation where FedAvg-ERM is the simple baseline (B). “-” means the method is not applicable in that context. Bold is for best and italics is for second best in each column. FedDG and FedADG are designed for image dataset and thus not applicable for CivilComments and Py150.
@@
FedAvg-ERM $0.5413$ $0.3929$ $0.3390$ $0.3331$ $\mathbf{0.6833}$ $\mathbf{0.6827}$ $0.6501$ $\mathit{0.6405}$
DG Adapted IRM $0.6408$ $\mathit{0.6153}$ $\mathbf{0.6093}$ $\mathit{0.6771}$ $0.6722$ $0.6379$ $0.6267$
Fish $\mathbf{0.6713}$ $\mathbf{0.6435}$ $\mathit{0.6086}$ $0.6634$ $0.6731$ $0.6464$ $0.6313$
MMD $\mathit{0.6520}$ $0.6121$ $\mathbf{0.6093}$ $0.6558$ $0.6731$ $0.6577$ $0.6565$
DeepCoral $0.5847$ $0.6121$ $\mathbf{0.6093}$ $0.6558$ $0.6723$ $0.6384$ $0.6236$
GroupDRO $0.6383$ $0.6028$ $0.4954$ $0.5127$ $\mathit{0.6824}$ $0.6421$ $0.6391$
FL FedProx - $0.4762$ $0.3635$ $\mathit{0.3409}$ $0.6730$ $\mathbf{0.6614}$ $\mathbf{0.6573}$
Scaffold - $0.4932$ $0.4930$ $\mathbf{0.4208}$ $0.6781$ $\mathit{0.6579}$ $0.6364$
FDG FedSR - 0.3603 0.3389 0.3188 - 0.0915 0.0860 0.0587
FedGMA - 0.3341 0.3157 0.3195 - 0.6628 0.3737 0.3576
Remark 4.2.
Most centralized DG methods degrade in the FL setting. For image datasets, the DG methods adapted to the FL setting (except for GroupDRO) show significant degradation in performance compared to their centralized counterparts as can be seen when comparing the $C=1$ column to the $C>1$ columns in Table 3. Further, degradation can be seen in PACS when moving from the homogeneous client setting ($\lambda=1$) to the heterogeneous client setting ($\lambda=0.1$).
Remark 4.3.
FL methods tackling client heterogeneity perform surprisingly well on the Federated DG task. FedProx and Scaffold, which were designed for client heterogeneity but not specifically for the DG task, perform quite well in the Federated DG setting and even perform the best for IWildCam and Py150. Thus, we suggest including these methods in future Federated DG evaluations and suggest that future works could focus on combining the strengths of these methods and DG methods.
Remark 4.4.
The performance of real-world data significantly degrades as $\lambda$ decreases. This can be seen from IWildCam and Py150. While it is challenging and expensive to run models for IWildCam and Py150, they show the largest differences between methods and demonstrates the real-world challenge of Federated DG. We suggest including IWildCam and Py150 in most future DG evaluations given their unique nature across datasets.
Remark 4.5.
A diversity of evaluation datasets is important for holistically evaluating new methods for Federated DG. This conclusion is inspired from the following two observations. 1) When comparing FL methods, we notice opposite trends between PACS and all other datasets (IWildCam, CivilComments, Py150) when increasing client heterogeneity based on $\lambda$. For PACS, the FL methods (except for FedADG and FedSR) surprisingly seem to improve when increasing client heterogeneity (i.e., $\lambda=0$). However, for other datasets, the expected trend exists such that increasing client heterogeneity produces worse performance. From this, we recommend that using PACS alone might be a misleading or at least incomplete evaluation of Federated DG methods. 2) Centralized DG methods adapted to FL perform the best on CivilComments. This unexpected result suggests that DG methods in the centralized regime may be able to better accommodate subpopulation shift, which is the special kind of shift that is exhibited in CivilComments. Both of these observations emphasize that a diversity of evaluation datasets is important for holistically evaluating new methods for Federated DG tasks.
Additional DG challenges from FL.
For further understanding, we explore some additional questions on the PACS dataset because it is the most common and computationally feasible to train many different models. Specifically, we explore how the number of clients, amount of communication (i.e., the number of server aggregations) in the federated setup, and client heterogeneity affects the performance of various methods. The figures and detailed analysis are provided in the Appendix but we highlight two remarks here.
Remark 4.6.
The number of clients $C$ strongly affects on the overall performance. The DG performance drops from $90\%$ to $50\%$ or even 10% when varying $C$ from $1$ to $200$. We strongly recommend future Federated DG evaluations to consider larger number of clients such as $50$ to $100$ rather than only a very small number of clients. It indicates that there exist significant unresolved challenges in the domain of federated domain generalization when dealing with a large number of clients, which have not been adequately explored. In particular, FedADG and FedSR seem to be sensitive to the number of clients because they perform poorly when $C\geq 10$, while in the original papers they were only evaluated on a few clients $C=3$ (and we reproduce the good results with such $C$ in the Appendix).
Remark 4.7.
The number of communications does not monotonically affect DG performance. We observe an interesting implicit regularization phenomena in the FL context: different methods achieve their best performance when the communication rounds is relatively low. In particular, for PACS, this communication round is $10$ (while fixing the total amount of computations). This is surprising as in contrast for in-domain task, FL suggests more communications leads to better performance (Stich, 2018, Theorem 5.1). Further theoretical investigation of the dependence of DG accuracy on the communications rounds, and its possible relation to implicit regularization via early stopping is an interesting area for future work.
5 Conclusion and Discussion
We first build a systematic benchmark methodology for evaluating Federated DG that includes novel methodologies for splitting the data across clients and evaluation of dataset difficulty. We then evaluate $13$ representative methods from three relevant lines of research. Our evaluation shows that Federated DG is still unsolved and significant gaps remain between centralized DG performance and Federated DG performance. Therefore, Federated DG is ripe for additional research. Based on our evaluation and observation, here are some recommendations and suggestions for future work in Federated DG.
Recommendations for future evaluations of Federated DG.
  1. 1.
    Stronger Baselines - FedAvg should always be included because it is a strong baseline (Remark 4.1). FL methods designed for client heterogeneity (though not necessarily DG) should also be included given their strong performance (Remark 4.3).
  2. 2.
    Realistic Datasets - Federated DG methods should be evaluated on more realistic datasets. In particular, FEMNIST and PACS behave quite differently with respect to client heterogeneity than the more realistic WILDS datasets (Remark 4.4). Thus, we recommend including both IWildCam and Py150 in future evaluations of Federated DG. CivilComments may be useful for evaluating realistic subpopulation shift (Remark 4.5).
  3. 3.
    Large Number of Heterogeneous Clients - Evaluations should include scenarios with a large number of clients with varying degrees of heterogeneity. A large number of clients poses unique challenges for most methods (Remark 4.6). Additionally, the heterogeneity of clients is more realistic and plays a significant role in evaluation (Remark 4.4).
Suggestions for future work in Federated DG.
  1. 1.
    Handle Domain Separation Case - The domain separation scenario ($\lambda=0$) limits the exchange of information between domains, i.e., a client may only have data from a single domain. Centralized approaches cannot be adapted to this setting and current methods still struggle under this realistic client heterogeneity setting.
  2. 2.
    Increase Convergence Rate - In most benchmark experiments (except for PACS, which is an easy dataset), we have observed slow convergence, particularly on challenging real-world datasets such as IWildCam and Py150. Moreover, we have noticed even slower convergence when client heterogeneity is high (i.e., when $\lambda$ is small). In addition, we have found that federated methods designed to improve convergence actually performed quite well even though they were not designed for the DG task (Remark 4.3). Therefore, it is crucial for future methods to improve the convergence rate of Federated DG approaches.
  3. 3.
    Investigate the Effect of Communication Frequency - Because increased communication frequency may actually hurt DG performance (Remark 4.7), further investigation into the regularization effect of early stopping and infrequent communications in FedAvg may yield important insights.
  4. 4.
    Understand the Effect of the Number of Clients - Given that the performance of some methods degrade very quickly when increasing the number of clients beyond a few (Remark 4.6), the community could benefit from an improved theoretic and empirical understanding of how the number of clients affects Federated DG in terms of convergence, computational requirements, sample complexity, and DG performance.
We hope this work provides a better foundation for future work in Federated DG and accelerates research progress.
Acknowledgements
Appendix
[…]
Reproducibility Statement
Code for reproduce the results is available at the following link:
https://github.com/inouye-lab/FedDG_Benchmark.
We include detailed documentation in using our code to reproduce the results throughout the paper. We also provide documentation in adding new algorithm’s DG evaluation in the FL context.
Appendix A Datasets and Difficulty Metric
A.1 Dataset Introduction
In this section, we introduce the datasets we used in our experiments, and the split method we used to build heterogeneous datasets in the training and testing phase as well as the heterogeneous local training datasets among clients in the FL.
FEMNIST
PACS
IWildCam
It is a real-world image classification dataset based on wild animal camera traps around the world, where each camera represents a domain. It contains 243 training domains, 32 validation and 48 test domains. Usually people cares about rare speicies, thus we utilize the macro F1 score as the evaluation metric instead of standard accuracy, as recommended in the original dataset’s reference Koh et al. [2021]. The $R_{\text{DG}}=0.449$ makes it a very challenging dataset for domain generalization.
CivilComments
It is a real-world binary classification text-based dataset formed from the comments from different demographic groups of people, containing $8$ demographic group. The goal is to judge whether the text is malicious or not.
Py150
A.2 Dataset Split Setup
For each dataset, we first split the dataset into $5$ categories, namely training dataset, in-domain validation dataset, in-domain test dataset, held-out validation dataset and test domain dataset. For FEMNIST and PACS dataset, we use Cartoon and Sketch as training domain, Art-painting as held-out domain, Painting as test domain. For training domain, we split $10\%,10\%$ of the total training domain datasets as in-domain validation dataset and in-domain test dataset respectively. For IWildCam, CivilComments and Py150, we directly apply the Wilds official splits.
Appendix B Benchmark Experimental Setting
B.1 Model Structure
In this benchmark, for image-based datasets, we use ResNet-50 He et al. [2016] dataset. For CivilComments and Py150 dataset, we choose DistilBERT Sanh et al. [2020] and CodeGPT Radford et al. [2019] respectively as recommended by Wilds.
B.2 Model Selection
B.3 Early Stopping
We conduct early stopping using the held-out validation dataset in our evaluation. For each dataset and method, We first run certain communication rounds, then we select the model parameters which achieves the best performance on the validation set. We report the held-out validation dataset in the main paper, and we report the results using the in-domain validation set in Appendix D.
B.4 Hyperparameters
In this section, we present the hyperparameters selected for the evaluation. We opted to communicate per epoch for all experiments. For PACS, we run totally $50$ communication rounds. For IWildCam, we run $50$ communication rounds. For CivilComments, we run $10$ communication rounds. For Py150, we run $10$ communication rounds, and for FEMNIST, we run $40$ communication rounds. Please refer to the Table 5 to review other hyperparameters.
Table 5: Hyperparameters used in the evaluation. c represents centralized learning regime, and f represents the federated learning regime
PACS IWildCam CivilComments Py150 FEMNIST
c f c f c f c f c f
Fed Avg Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
IRM Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Penalty weight $100$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $100$
Penalty anneal iters $500$ $500$ $500$ $500$ $500$ $500$ $500$ $500$ $500$ $500$
Fish Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
meta learning rate $0.5$ $0.5$ $0.5$ $0.5$ $0.5$ $0.5$ $0.5$ $0.5$ $0.5$ $0.5$
Mixup Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
$\alpha$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$
MMD Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Penalty Weight $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$
Deep Coral Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Penalty Weight $1$ $1$ $10$ $10$ $1$ $1$ $1$ $1$ $1$ $1$
Group DRO Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
$\eta$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$
Fed Prox Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
$\mu$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$
Scaf fold Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
FedDG Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
$\lambda$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$
FedADG Classifier optimizer SGD SGD SGD SGD SGD SGD SGD SGD SGD SGDs
Classifier learning rate $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$
Gen optimizer SGD SGD SGD SGD SGD SGD SGD SGD SGD SGD
Gen learning rate $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$
Disc optimizer SGD SGD SGD SGD SGD SGD SGD SGD SGD SGD
Disc learning rate $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$
$\alpha$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$
FedSR Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
$\ell_{2}$ regularizer $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.001$ $0.001$ $0.01$ $0.01$
cmi regularizer $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$
FedGMA Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Server step size $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$
Mask threshold $0.4$ $0.4$ $0.4$ $0.4$ $0.4$ $0.4$ $0.4$ $0.4$ $0.4$ $0.4$
Appendix C Additional FL-specific challenges for domain generalization
As mentioned in Section 4, we also include some deeper exploration over the effect of number of clients and communication frequency, which are unique to the FL regime.
i) Massive number of clients: In this experiment, we explore the performance of different algorithms when the number of clients $K$ increases on PACS. We fix the communication rounds $50$ and the local number of epoch is $1$ (synchronizing the models every epoch). Fig. 2 plots the held-out DG test accuracy versus number of clients for different levels of data heterogeneity. The following comments are in order: given communication budget, 1) current domain generalization methods all degrade a lot after $C\geq 10$, while the performance ERM and FedDG maintain relatively unchanged as the clients number increases given communication budget. FedADG and FedSR are are sensitive to the clients number, and they both fail after $C\geq 20.$ 2) Even in the simplest homogeneous setting $\lambda=1,$ where each local client has i.i.d training data, current domain generalization methods IRM, FISH, Mixup, MMD, Coral, GroupDRO work poorly in the existence of large clients number, this means new methods are needed for DG in FL context when data are stored among massive number of clients.
Figure 2: PACS: Held-out DG test accuracy versus number of clients.
Appendix D Supplementary Results
In the main paper, we provide experiments results using held-out validation early stopping. Here we report the results using in-domain validation set. We also report the convergence curve for each methods on each dataset for reference. We observe that under most of cases, the held-out validation gives us a better model. Thus, we recommend using held-out validation set to perform early stopping.
More results on PACS and IWildCam dataset.
Here we both report the convergence curve using the DG accuracy of each method on PACS Fig. 4 and IWildCam Fig. 5 and the DG accuracy on PACS and IWildCam Table 6 using the in-domain validation. From the PACS convergence curve Fig. 4, we observe that with lower $\lambda$, the model is harder to converge. FedAvg ang FedDG converge faster than all other methods while FedADG and FedSR does not converge. From IWildCam convergence curve Fig. 5, we observe that all methods are struggling to converge. This is due to the main challenge first come from the $R_{\text{FL}}$ here. We thus observe that Federated methods dealing with heterogeneity clients achieves best performance. Especially when $\lambda$ is low. It is worth noticing that this does not mean DG in FL is solved, none of the methods even achieve the centralized ERM performance.
Table 6: Test accuracy on PACS and IWildCam dataset with in-domain validation where FedAvg-ERM is the simple baseline (B). “-” means the method is not applicable in that context. Bold is for best and italics is for second best in each column.
PACS ($D=2$) IWildCam ($D=243$)
$C=1$ $C=100$ (FL) $C=1$ $C=243$ (FL)

### v2

@@ Appendix F Supplementary Results
Table 8: Test accuracy on PACS and IWildCam dataset with held-out-domain validation where FedAvg-ERM is the simple baseline (B). “-” means the method is not applicable in that context. Bold is for best and italics is for second best in each column. We report the standard deviation among 3 runs.
@@
FedAvg-ERM $\bf 0.943\pm 0.012$ $\bf 0.948\pm 0.020$ $\mathbf{0.955\pm 0.016}$ $\mathbf{0.954\pm 0.018}$ $\mathit{0.343\pm 0.003}$ $\mathbf{0.298\pm 0.002}$ $0.245\pm 0.002$ $\mathbf{0.196\pm 0.010}$
DG Adapted IRM $0.918\pm 0.007$ $0.920\pm 0.041$ $0.856\pm 0.044$ $0.321\pm 0.004$ $0.196\pm 0.006$ $0.179\pm 0.005$
Fish $\mathit{0.936\pm 0.021}$ $0.653\pm 0.127$ $0.345\pm 0.107$ $\mathbf{0.354\pm 0.001}$ $0.183\pm 0.004$ $0.154\pm 0.005$
Mixup $0.918\pm 0.006$ $0.886\pm 0.052$ $0.826\pm 0.030$ $0.331\pm 0.007$ $0.086\pm 0.010$ $0.072\pm 0.007$
MMD $0.928\pm 0.022$ $0.921\pm 0.041$ $0.855\pm 0.044$ $0.324\pm 0.002$ $0.185\pm 0.005$ $0.163\pm 0.007$
DeepCoral $0.928\pm 0.007$ $0.920\pm 0.041$ $0.856\pm 0.043$ $0.329\pm 0.008$ $0.185\pm 0.005$ $0.163\pm 0.007$
GroupDRO $0.928\pm 0.007$ $\mathit{0.944\pm 0.025}$ $0.945\pm 0.015$ $0.211\pm 0.002$ $0.134\pm 0.005$ $0.198\pm 0.006$
FL FedProx - $0.896\pm 0.025$ $0.891\pm 0.025$ $0.896\pm 0.024$ $0.251\pm 0.003$ $0.204\pm 0.001$ $0.167\pm 0.003$
Scaffold - $0.896\pm 0.025$ $0.892\pm 0.024$ $0.896\pm 0.024$ $\mathit{0.281\pm 0.002}$ $\mathbf{0.255\pm 0.004}$ $\mathit{0.178\pm 0.008}$
AFL - $0.931\pm 0.026$ $0.915\pm 0.040$ $0.911\pm 0.041$ $0.256\pm 0.010$ $0.162\pm 0.008$ $0.034\pm 0.001$
FDG FedDG - $0.933\pm 0.020$ $\mathit{0.947\pm 0.018}$ $\mathit{0.943\pm 0.023}$ $0.274\pm 0.001$ $0.235\pm 0.002$ $0.167\pm 0.004$
FedADG - $0.943\pm 0.011$ $0.942\pm 0.001$ $0.935\pm 0.011$ $0.259\pm 0.012$ $\mathit{0.251\pm 0.010}$ $0.164\pm 0.001$
FedSR - $0.640\pm 0.158$ $0.548\pm 0.086$ $0.537\pm 0.092$ $0.177\pm 0.005$ $0.141\pm 0.005$ $0.086\pm 0.003$
FedGMA - $0.750\pm 0.050$ $0.730\pm 0.087$ $0.724\pm 0.099$ $0.223\pm 0.001$ $0.151\pm 0.002$ $0.091\pm 0.003$
Table 9: Test accuracy on CelebA and Camelyon17 dataset with held-out-domain validation where FedAvg-ERM is the simple baseline (B). “-” means the method is not applicable in that context. Bold is for best and italics is for second best in each column. We report the standard deviation among 3 runs.
CelebA ($D=2$) Camelyon17 ($D=3$)
$C=1$ $C=100$ (FL) $C=1$ $C=100$ (FL)
B
FedAvg-ERM $0.769\pm 0.035$ $0.606\pm 0.019$ $0.517\pm 0.039$ $0.446\pm 0.018$ $0.903\pm 0.009$ $\mathit{0.949\pm 0.007}$ $\mathit{0.950\pm 0.004}$ $\bf 0.948\pm 0.004$
DG Adapted IRM $\mathbf{0.891\pm 0.012}$ $0.706\pm 0.048$ $0.737\pm 0.013$ $0.912\pm 0.056$ $0.948\pm 0.001$ $0.952\pm 0.002$
Fish $\mathit{0.883\pm 0.010}$ $0.144\pm 0.158$ $0.072\pm 0.054$ $0.922\pm 0.006$ $0.881\pm 0.006$ $0.878\pm 0.008$
Mixup $0.288\pm 0.498$ $0.129\pm 0.217$ $0.126\pm 0.218$ $\mathit{0.940\pm 0.011}$ $\mathbf{0.949\pm 0.003}$ $0.947\pm 0.006$
MMD $0.835\pm 0.047$ $0.809\pm 0.027$ $0.766\pm 0.030$ $0.922\pm 0.027$ $0.946\pm 0.004$ $0.947\pm 0.006$
DeepCoral $0.852\pm 0.036$ $\mathbf{0.813\pm 0.017}$ $\mathbf{0.782\pm 0.022}$ $\mathbf{0.940\pm 0.007}$ $0.948\pm 0.004$ $0.949\pm 0.003$
GroupDRO $0.869\pm 0.034$ $\mathit{0.811\pm 0.028}$ $0.761\pm 0.056$ $0.919\pm 0.027$ $0.947\pm 0.006$ $\mathbf{0.953\pm 0.001}$
FL FedProx - $0.126\pm 0.237$ $0.121\pm 0.245$ $000.000\pm 0$ $0$ $0.939\pm 0.002$ $0.935\pm 0.001$ $\mathit{0.937\pm 0.007}$
Scaffold - $0.728\pm 0.010$ $\mathit{0.776\pm 0.016}$ $\bf 0.800\pm 0.017$ $0.942\pm 0.001$ $0.929\pm 0.003$ $0.932\pm 0.001$
AFL - $0.811\pm 0.024$ $0.834\pm 0.005$ $0.828\pm 0.005$ $0.942\pm 0.011$ $0.947\pm 0.006$ $0.932\pm 0.008$
FDG FedDG - $0.561\pm 0.184$ $0.531\pm 0.112$ $0.464\pm 0.201$ $0.863\pm 0.012$ $0.856\pm 0.003$ $0.869\pm 0.016$
FedADG - $0.674\pm 0.003$ $0.669\pm 0.008$ $\mathit{0.600\pm 0.009}$ $0.936\pm 0.001$ $0.933\pm 0.004$ $0.926\pm 0.002$
FedSR - $0.381\pm 0.018$ $0.363\pm 0.041$ $0.383\pm 0.015$ $0.934\pm 0.010$ $0.917\pm 0.009$ $0.925\pm 0.003$
FedGMA - $0.620\pm 0.012$ $0.615\pm 0.018$ $0.487\pm 0.008$ $0.901\pm 0.006$ $0.854\pm 0.006$ $0.815\pm 0.025$
Table 10: Test accuracy on CivilComments and Py150 datasets with held-out-domain validation where FedAvg-ERM is the simple baseline (B). “-” means the method is not applicable in that context. Bold is for best and italics is for second best in each column. FedDG and FedADG are designed for image dataset and thus not applicable for CivilComments and Py150. We report the standard deviation among 3 runs.
@@
FedAvg-ERM $0.541\pm 0.003$ $0.359\pm 0.028$ $0.347\pm 0.018$ $0.325\pm 0.009$ $\mathbf{0.683}\pm 0.003$ $\mathbf{0.683}\pm 0.000$ $\mathit{0.650\pm 0.000}$ $\mathit{0.641}\pm 0.000$
DG Adapted IRM $0.641\pm 0.001$ $\mathit{0.587\pm 0.018}$ $0.534\pm 0.038$ $\mathit{0.677}\pm 0.000$ $\mathit{0.655\pm 0.001}$ $\mathbf{0.653\pm 0.000}$ $0.627\pm 0.001$
Fish $\mathbf{0.671}\pm 0.000$ $0.424\pm 0.158$ $0.340\pm 0.173$ $0.663\pm 0.000$ $0.650\pm 0.000$ $0.648\pm 0.001$ $\mathbf{0.642\pm 0.001}$
MMD $\mathit{0.652}\pm 0.000$ $\mathbf{0.634\pm 0.012}$ $\mathbf{0.613\pm 0.014}$ $0.656\pm 0.001$ $0.627\pm 0.000$ $0.629\pm 0.000$ $0.610\pm 0.000$
DeepCoral $0.585\pm 0.000$ $0.515\pm 0.061$ $0.463\pm 0.080$ $0.656\pm 0.003$ $0.651\pm 0.000$ $0.650\pm 0.001$ $0.639\pm 0.000$
GroupDRO $0.638\pm 0.003$ $0.475\pm 0.014$ $\mathit{0.474\pm 0.003}$ $0.513\pm 0.001$ $0.589\pm 0.000$ $0.603\pm 0.003$ $0.607\pm 0.002$
FL FedProx - $0.181\pm 0.032$ $0.173\pm 0.033$ $0.170\pm 0.036$ $0.638\pm 0.000$ $0.633\pm 0.000$ $0.612\pm 0.000$
Scaffold - $0.389\pm 0.015$ $0.376\pm 0.018$ $\mathit{0.332\pm 0.014}$ $0.642\pm 0.000$ $0.638\pm 0.000$ $0.617\pm 0.000$
AFL - $0.552\pm 0.021$ $0.474\pm 0.014$ $\mathbf{0.435\pm 0.019}$ $0.486\pm 0.000$ $0.485\pm 0.003$ $0.467\pm 0.001$
FDG FedSR - $0.360\pm 0.003$ $0.339\pm 0.003$ $0.319\pm 0.003$ $0.533\pm 0.002$ $0.526\pm 0.001$ $0.445\pm 0.003$
FedGMA - $0.209\pm 0.028$ $0.202\pm 0.024$ $0.200\pm 0.022$ $0.620\pm 0.000$ $0.613\pm 0.000$ $0.600\pm 0.000$
Table 11: Test accuracy on FEMNIST dataset with held-out validation where FedAvg-ERM is the simple baseline (B). “-” means the method is not applicable in that context. Bold is for best and italics is for second best in each column.
FEMNIST (Held-Out-Domain)
$C=1$ $C=100$ (FL)
(centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$